OpenStudy (anonymous):

How many 3x3 matrices 'A' can you find such that x x+y A y = x-y z 0 for all choices of x, y and z?

OpenStudy (anonymous):

\[A \left(\begin{matrix}x \\ y\ \\z\end{matrix}\right) =\left(\begin{matrix}x+y \\ x-y\\0\end{matrix}\right)\]

OpenStudy (loser66):

just let a11, a12, a13..... as entries of A then you have a11x +a12 y +a13 z = x +y

OpenStudy (rational):

I can only think of one so far \[\large \left(\begin{matrix} 1&1&0\\1&-1&0\\0&0&0\end{matrix} \right) \left(\begin{matrix}x \\ y\ \\z\end{matrix}\right) =\left(\begin{matrix}x+y \\ x-y\\0\end{matrix}\right)\]

OpenStudy (loser66):

make a comparison to get a11 = 1, a12 = 1 a13 =0

OpenStudy (anonymous):

@rational , your answer is correct, but how did you arrive to that. @Loser66 i tried, but there's too many unknowns

OpenStudy (loser66):

@rational the last line can't be 0, it makes A is not invertible.

OpenStudy (rational):

it has to be 0 for arbitrary x y z

OpenStudy (rational):

I took linear combinations of rows of (x,y,z) since the matrix A was left multiplied to the vector (x,y,z) @mickey4691

OpenStudy (loser66):

ok, let me show my work for the last one, we have a31 x +a32y +a33z =0, one of a can be 0 and others is 1 and -1 the permutation of it is 6, so that we have 6 matrices

OpenStudy (anonymous):

but you are assuming that one of them is zero which would defy the purpose of proving? @Loser66 @rational what do you mean by linear combinations?

OpenStudy (rational):

that won't result in 0 for z component right ? because a31 = 0, a32=1,a33=-1 gives y-z, not 0

OpenStudy (rational):

left multiplying a vector by a matrix is same as taking linear combinations of rows of vector

OpenStudy (loser66):

OH yeah!!!! My bad. I am sooooo sorry.

OpenStudy (rational):

\[\large \left(\begin{matrix} a&b&c \end{matrix} \right) \left(\begin{matrix}x \\ y\ \\z\end{matrix}\right) =\left(\begin{matrix}ax +by+cz\end{matrix}\right)\]

OpenStudy (rational):

we can interpret this multiplication as taking linear combinations of "rows" of second matrix: take "a" of first row take "b" of second row take "c" of third row and add them

OpenStudy (rational):

So if you want \(ax-by\) as result, you would do : \[\large \left(\begin{matrix} a&-b&0 \end{matrix} \right) \left(\begin{matrix}x \\ y\ \\z\end{matrix}\right) =\left(\begin{matrix}ax -by\end{matrix}\right)\] right ?

OpenStudy (anonymous):

I see what you're doing there but this method is more of a inspection and comparison? correct?

OpenStudy (rational):

nope, this is one of the four ways of multiplying two matrices

OpenStudy (rational):

and the result of matrix multiplication is unique so we won't be missing out any solutions i hope

OpenStudy (anonymous):

hmm, i will attach a photo of my working. Give me a few. Thanks. I understand

OpenStudy (rational):

Okay...

OpenStudy (loser66):

yup

OpenStudy (rational):

Looks good !

OpenStudy (anonymous):

Thank you @rational and @Loser66

OpenStudy (rational):

yw :)